# NeuroTree: Hierarchical Functional Brain Pathway Decoding for Mental Health Disorders

### v2

@@ 11 Conclusion
This study proposed NeuroTree, a novel hierarchical brain tree framework that integrates k-hop AGE-GCN with neural ODEs and an attention-based contrastive loss to decode functional brain pathways in psychiatric disorders. NeuroTree effectively incorporates demographic factors to improve brain disease classification and enhances interpretability in predicting brain age while revealing distinct hierarchical connectivity patterns within the DMN, FPN, and VAN across different mental disorders. Additionally, these findings demonstrate that different subtypes of mental disorders exhibit varying trajectories of FC changes across age periods. These insights contribute to the neuroscientific understanding of functional alterations induced by mental disorders, potentially informing therapeutic interventions and pharmacological treatments.
[…]
12 Appendix: Proof Formula 7 for K-hop Connectivity
(K-hop Connectivity) Given the $k$-hop connectivity adjacency operator:
@@ 12 Appendix: Proof Formula 7 for K-hop Connectivity
$\big{[}\lambda A^{d}(t)+(1-\lambda)(A^{d}(t))^{T}\big{]}^{k}$ (26)
$=\sum_{i=0}^{k}\binom{k}{i}\lambda^{i}(1-\lambda)^{k-i}\big{(}A^{d}(t)\big{)}^{i}\big{(}(A^{d}(t))^{T}\big{)}^{k-i}.$
Proof 12.1.
@@
$\big{[}\lambda A^{d}(t)+(1-\lambda)(A^{d}(t))^{T}\big{]}^{n}=\sum_{i=0}^{n}$ $\binom{n}{i}\lambda^{i}(1-\lambda)^{n-i}$
×$\kern-19.91684pt\times\big{(}A^{d}(t)\big{)}^{i}\big{(}(A^{d}(t))^{T}\big{)}^{n-i}.$
@@
$\big{[}\lambda A^{d}(t)+(1-\lambda)(A^{d}(t))^{T}\big{]}^{n+1}$
$=$ $\big{[}\lambda A^{d}(t)+(1-\lambda)(A^{d}(t))^{T}\big{]}^{n}$
×$\kern-19.91684pt\times\big{[}\lambda A^{d}(t)+(1-\lambda)(A^{d}(t))^{T}\big{]}.$
@@
$\big{[}\lambda A^{d}(t)+(1-\lambda)(A^{d}(t))^{T}\big{]}^{n+1}$
$=\sum_{i=0}^{n+1}$ $\binom{n+1}{i}\lambda^{i}(1-\lambda)^{n+1-i}$
×$\kern-19.91684pt\times\big{(}A^{d}(t)\big{)}^{i}\big{(}(A^{d}(t))^{T}\big{)}^{n+1-i}.$
Thus, by induction, the formula holds for all $k$. And we have following corollary in 12.2
Corollary 12.2 (Asymmetric Property of $K$-hop Operator).
@@
Applying the binomial expansion from 12:
@@
$(\hat{A}_{k}(t))^{T}$ $=\Gamma^{T}\odot(A^{s})^{T}\odot\sum_{i=0}^{k}\binom{k}{i}\lambda^{i}(1-\lambda)^{k-i}$ (30)
×$\qquad\times((A^{d}(t))^{i})^{T}(((A^{d}(t))^{T})^{k-i})^{T}$
$=\Gamma\odot A^{s}\odot\sum_{i=0}^{k}\binom{k}{i}\lambda^{i}(1-\lambda)^{k-i}$
×$\qquad\times((A^{d}(t))^{T})^{i}(A^{d}(t))^{k-i}$
@@
Therefore, $\hat{A}_{k}(t)\neq(\hat{A}_{k}(t))^{T}$ unless all terms in the expansion are zero.
13 Appendix: Proof of Theorem  4.1
We prove that $\|\hat{A}_{k}(t)\|_{2}$ has an upper bound as $k$-hop approaches infinity:
@@
Using equation 35, we bound:
$\|[\lambda A^{d}(t)+(1-\lambda)(A^{d}(t))^{T}]\|_{2}$ $\leq\max(\lambda,1-\lambda)$ (38)
×$\qquad\kern-19.91684pt\times(\|A^{d}(t)\|_{2}+\|(A^{d}(t))^{T}\|_{2})$
@@
Taking the power of $k$ on both sides of equation 39: $\|\hat{A}_{k}(t)\|_{2}\leq\|\hat{A}(t)\|_{2}^{k}$ and the following inequality holds:
@@
Since $\max(\lambda,1-\lambda)<1$ for $\lambda\in(0,1)$, the series converges geometrically. This result indicates that $\|\hat{A}_{k}(t)\|$ converges to zero as $k\to\infty$, provided the spectral radius of $A$, denoted by $\lambda$, satisfies $\max(\lambda,1-\lambda)<1$. Consequently, the rate of decay of $\|\hat{A}_{k}(t)\|$ is governed by $\max(\lambda,1-\lambda)^{k}$, which demonstrates the dependence on the spectral properties of $A$ and ensures convergence for sufficiently small $\lambda$.
14 Appendix: Proof for Proposition I
To prove the proposition I, we express the neural ODE system as:
@@ 14 Appendix: Proof for Proposition I
Thus, Eq. 48 ensures that $F$ is continuous in $t$. According to Picard’s existence theorem, the $k$-hop ODE-GCN has a unique solution.
15 Appendix: Finding the Shortest Paths in a Weighted Brain Tree
@@
(Optimal Path Extremes) Let $\mathcal{W}(S)=\sum_{v\in P}S(v;\boldsymbol{\Theta})$ represent the node score contribution and $\mathcal{W}(C)=\sum_{e_{ij}\in E(P)}\mathcal{C}_{i,j}$ represent the connectivity strength contribution in the pruned graph $\mathcal{T}(G)$. For an arbitrary integer threshold $\lambda^{*}$, the optimal path $P^{*}$ satisfies the following interval constraint $\mathcal{W}(S)\leq\lambda^{*}\leq\mathcal{W}(C)$.
@@
In practice, the minimum interval can be determined using the bisection method within $[\alpha_{L},\alpha_{U}]$ using a bisection method such that $\mathcal{W}(S)\leq\mathcal{W}_{\alpha_{L}}(P^{*})\leq\alpha^{*}\leq\mathcal{W}_{\alpha_{U}}(P^{*})\leq\mathcal{W}(C)$ [40].

### v3

@@ 10 Conclusion
This study proposed NeuroTree, a novel hierarchical brain tree framework that integrates $k$-hop AGE-GCN with neural ODEs and an attention-based contrastive loss to decode functional brain pathways in psychiatric disorders. NeuroTree effectively incorporates demographic to improve brain disease classification and enhances interpretability in predicting brain age while revealing distinct hierarchical connectivity patterns within the DMN, FPN, and VN across different mental disorders. Additionally, these findings demonstrate that different subtypes of mental disorders exhibit varying trajectories of FC changes across age periods. These insights contribute to the neuroscientific understanding of functional alterations induced by mental disorders, potentially informing therapeutic interventions and pharmacological treatments.
Impact Statement
In this study, we examine the increasing societal impact of mental illness and addiction, highlighting the critical importance of understanding the underlying mechanisms of brain disorders. Traditional GNNs designed for fMRI data have proven effective in capturing network-based features; however, many of these approaches lack interpretability and fail to incorporate external demographic information within a deep learning framework. To address these limitations, we introduce NeuroTree, a novel framework that establishes a robust foundation for advancing neuroscience research, identifies influential brain network regions associated with disorders, and elucidates potential therapeutic targets for future drug development and intervention strategies in specific brain regions. In future work, our NeurTree framework can be extended to related research fields such as EEG and social networks.
Reproducibility
All of our experiments are reproducible, with results presented in Table 1 and Table 2. We will upload the preprocessed data to Google Drive and release our code on GitHub upon acceptance of the paper. Please note that, based on our experience, there may be some variations in the results shown in Table 1 and Table 2 due to differing GPU configurations. For accurate results, please refer to our released checkpoints.
Appendix A Appendix: Proof Formula 7 for K-hop Connectivity
Theorem A.1 (K-hop Connectivity).
Given the $k$-hop connectivity adjacency operator:
@@ Appendix A Appendix: Proof Formula 7 for K-hop Connectivity
$\big{[}\lambda A^{d}(t)+(1-\lambda)(A^{d}(t))^{T}\big{]}^{k}=\sum_{i=0}^{k}\binom{k}{i}\lambda^{i}(1-\lambda)^{k-i}\big{(}A^{d}(t)\big{)}^{i}\big{(}(A^{d}(t))^{T}\big{)}^{k-i}.$ (19)
Proof.
@@
$\big{[}\lambda A^{d}(t)+(1-\lambda)(A^{d}(t))^{T}\big{]}^{n}=\sum_{i=0}^{n}\binom{n}{i}\lambda^{i}(1-\lambda)^{n-i}\big{(}A^{d}(t)\big{)}^{i}\big{(}(A^{d}(t))^{T}\big{)}^{n-i}.$
@@
$\big{[}\lambda A^{d}(t)+(1-\lambda)(A^{d}(t))^{T}\big{]}^{n+1}=\big{[}\lambda A^{d}(t)+(1-\lambda)(A^{d}(t))^{T}\big{]}^{n}\big{[}\lambda A^{d}(t)+(1-\lambda)(A^{d}(t))^{T}\big{]}.$
@@
$\big{[}\lambda A^{d}(t)+(1-\lambda)(A^{d}(t))^{T}\big{]}^{n+1}=\sum_{i=0}^{n+1}\binom{n+1}{i}\lambda^{i}(1-\lambda)^{n+1-i}\big{(}A^{d}(t)\big{)}^{i}\big{(}(A^{d}(t))^{T}\big{)}^{n+1-i}.$
Thus, by induction, the formula holds for all $k$. And we have following corollary in A.2
Corollary A.2 (Asymmetric Property of $K$-hop Operator).
@@
Applying the binomial expansion from A.1:
@@
$(\hat{A}_{k}(t))^{T}$ $=\Gamma^{T}\odot(A^{s})^{T}\odot\sum_{i=0}^{k}\binom{k}{i}\lambda^{i}(1-\lambda)^{k-i}((A^{d}(t))^{i})^{T}(((A^{d}(t))^{T})^{k-i})^{T}$ (23)
$=\Gamma\odot A^{s}\odot\sum_{i=0}^{k}\binom{k}{i}\lambda^{i}(1-\lambda)^{k-i}((A^{d}(t))^{T})^{i}(A^{d}(t))^{k-i}$
@@
Therefore, $\hat{A}_{k}(t)\neq(\hat{A}_{k}(t))^{T}$ unless all terms in the expansion are zero. ∎
Appendix B Appendix: Proof of Theorem  4.2
We prove the $\|\hat{A}_{k}(t)\|_{2}$ has an upper bound as $k$-hop approaches infinity:
@@
Using Eq. 27, we further bound:
$\|[\lambda A^{d}(t)+(1-\lambda)(A^{d}(t))^{T}]\|_{2}\leq\max(\lambda,1-\lambda)(\|A^{d}(t)\|_{2}+\|(A^{d}(t))^{T}\|_{2})$ (30)
@@
Taking the power of $k$ on both sides of Eq. 31, we have $\|\hat{A}_{k}(t)\|_{2}\leq\|\hat{A}(t)\|_{2}^{k}$ and the following inequality holds:
@@
Since $\max(\lambda,1-\lambda)<1$ for $\lambda\in(0,1)$, the series converges geometrically. This result indicates that $\|\hat{A}_{k}(t)\|$ converges to zero as $k\to\infty$, provided the spectral radius of $A$, denoted by $\lambda$, satisfies $\max(\lambda,1-\lambda)<1$. Consequently, the rate of decay of $\|\hat{A}_{k}(t)\|$ is governed by $\max(\lambda,1-\lambda)^{k}$, which demonstrates the dependence on the spectral properties of $A$ and ensures convergence for sufficiently small $\lambda$. ∎
Appendix C Appendix: Proof for Proposition 4.3
To prove the proposition 4.3, we express the neural ODE system as:
@@
Thus, Eq. 40 ensures that $F$ is continuous in $t$. According to Picard’s existence theorem, the $k$-hop ODE-GCN has a unique solution. ∎
Appendix D Appendix: Finding the Shortest Paths in a Weighted Brain Tree
Theorem D.1 (Optimal Path Extremes).
@@ Appendix D Appendix: Finding the Shortest Paths in a Weighted Brain Tree
Let $\mathcal{W}(S)=\sum_{v\in P}S(v;\boldsymbol{\Theta})$ represent the node score contribution and $\mathcal{W}(C)=\sum_{e_{ij}\in E(P)}\mathcal{C}_{i,j}$ represent the connectivity strength contribution in the pruned graph $\mathcal{T}(G)$. For an arbitrary integer threshold $\lambda^{*}$, the optimal path $P^{*}$ satisfies the following interval constraint $\mathcal{W}(S)\leq\lambda^{*}\leq\mathcal{W}(C)$.
@@
In practice, the minimum interval can be determined using the bisection method within $[\alpha_{L},\alpha_{U}]$ using a bisection method such that $\mathcal{W}(S)\leq\mathcal{W}_{\alpha_{L}}(P^{*})\leq\alpha^{*}\leq\mathcal{W}_{\alpha_{U}}(P^{*})\leq\mathcal{W}(C)$ (Xue, 2000).
Appendix E Prediction of FC Alterations in Brain Age Across Different Age Groups
This section presents our application of NeuroTree for predicting mental disorders and examining their impact on brain FC changes across various age cohorts. We utilized AGE-GCN embeddings to reconstruct FC values for different age groups. Subsequently, we analyzed FC values for various age groups, as illustrated in Fig. LABEL:fig:age_comparision. The results presented in Fig. LABEL:fig:age_comparision (a) and (b) reveal that significant within-group differences were observed exclusively among cannabis users in the 18–25 age group. We observe significant changes in FC values associated with cannabis use disorder and schizophrenia across various age groups, indicating age-dependent variations in FC patterns among individuals with these psychiatric disorders.
Appendix F High-Order Tree Path Influence
In further examining the characterization of fMRI in the tree path, we investigated the contribution of FC strength through Eq. 16 across two distinct subtypes of fMRI representations, as illustrated in Fig. 7 Our observations revealed that cannabis-related pathways exhibited relatively higher average FC weights compared to the COBRE dataset. Notably, both the addiction and schizophrenia groups demonstrated significant decreases following an increase in $\alpha$, while the HC cohort maintained relatively stable patterns. These findings provide compelling evidence that higher-order brain tree paths effectively capture the differential contributions of FC strength arising from alterations related to addiction or schizophrenia.
Figure 7: Average FC value weights versus $\alpha$ in high-order tree paths, comparing (a) HC vs cannabis users and (b) HC versus schizophrenia patients. The results indicate reduced connectivity patterns with increasing $\alpha$ in both clinical groups compared to the control group.
Appendix G More Related Work
Schizophrenia (SCZ) and substance abuse (SA) both profoundly impact brain structure and function, with a particularly high comorbidity rate observed among younger populations. Research indicates that SCZ patients frequently exhibit functional abnormalities in critical brain regions, including the DMN, central executive network (CEN), and visual pathways (Ma et al., 2023). However, findings remain inconsistent regarding whether comorbid SA history in SCZ patients exacerbates these neural impairments. (Wojtalik & Barch, 2014) conducted fMRI scans during working memory tasks and discovered significantly enhanced activation in frontal-parietal-thalamic regions among control subjects with SA history. Interestingly, SCZ patients showed no significant activation differences regardless of SA history, suggesting that pre-existing neural deficits in SCZ may obscure the additional effects of SA (Wojtalik & Barch, 2014). In addition, (Passiatore et al., 2023) further demonstrated that FC alterations associated with SCZ risk emerge during adolescence, particularly in cerebellum-occipitoparietal and prefrontal-sensorimotor regions, strongly correlating with polygenic risk scores (PRS).
Beyond network abnormalities, SCZ and SA patients commonly exhibit accelerated brain aging, where their brain-predicted age significantly exceeds their chronological age, indicating heightened functional deterioration of the brain—a phenomenon with critical implications for understanding cognitive decline and disease risk (Cole & Franke, 2017; Lombardi et al., 2021). With the emergence of GNN applications to fMRI data, researchers can now model and analyze brain regions as nodes and functional connections as edges in graph structures. The Pooling Regularized GNN (PR-GNN) proposed by (Li et al., 2020). Applying this methodology to SCZ with comorbid SA may simultaneously facilitate classification and biomarker identification, while incorporating demographic variables such as age, gender, and SA history enables individualized modeling. In summary, SCZ and SA have age-related and highly intertwined effects on brain functional networks. GNN methodologies provide technical tools with both explanatory and predictive capabilities, offering promise for deepening our understanding of the neural mechanisms underlying psychiatric disorders and their potential clinical applications.
Appendix H Experiments Detail
Experiments Setting
Baseline Models.
The classical methods for constructing brain networks include Pearson correlation and $k$-nearest neighbor approaches combined with GCN, such as:These methods are as follows:1) $Pearson$ GCN: This model constructs an fMRI graph using Pearson correlation coefficients to represent relationships between ROIs. 2) $k$-NN GCN: This model builds brain networks by connecting k-nearest neighbors based on similar input features within brain regions in the graph. 3) GAT (Veličković et al., 2017): This model extends the GCN framework by incorporating an attention mechanism that enhances the aggregation and learning of features from neighboring nodes. 4) BrainGNN (Li et al., 2021): BrainGNN is an interpretable GNN framework designed for analyzing fMRI data. This framework effectively identifies and characterizes significant brain regions and community patterns associated with specific neurological conditions. 5) BrainUSL (Zhang et al., 2023): BrainUSL learns the graph structure directly from BOLD signals while incorporating sparsity constraints and contrastive learning to capture meaningful connections between brain regions. Finally, we compare NeuroTree with other methods that incorporate higher-order node information, specifically MixHop (Abu-El-Haija et al., 2019) and graph path-based models such as GPC-GCN (Li et al., 2022b) and PathNN (Michel et al., 2023).
Evaluation Metrics
In our fMRI graph classification experiments (Section 6.2), we evaluate NeuroTree using Area Under the Curve (AUC), Accuracy (Acc.), Precision (Prec.), and Recall (Rec.). In brain age prediction (Section 8), we evaluate NeuroTree predicted age and chronological age using MSE metric.
Parameters Setting
In this study, two datasets were trained with a batch size of 16 for 100 epochs using a learning rate of 0.001.
Table 3: Training Parameters.
Notation Meaning Value
$\rho$ Scale parameter $0.5$
$T$ Number of time segment 2
$\lambda$ $k$-hop connectivity balance between dynamic FC matrices $A^{d}(t)$ (0,1)
$\beta$ Learnable age-modulated parameter (0,1)
$\Gamma$ Learnable graph weight parameter -
$W_{k}^{(l)}$ Learnable weight matrix in $l$th layer -
Additional Ablation Study
In this subsection, we conduct ablation experiments on the performance of brain networks classification with respect to two parameters $\lambda$ and $\Gamma$ on two mental disorders dataset. The experimental results demonstrate that for Cannabis, the model achieves optimal accuracy (0.8) and AUC value (0.87) at $\lambda=0.2$, with performance decreasing as $\lambda$ increases to 0.4, followed by a recovery at $\lambda=0.6$. In contrast, COBRE exhibits a different parameter sensitivity pattern, with accuracy showing a decreasing-then-increasing trend as $\lambda$ increases from 0.2 to 0.8, ultimately reaching peak accuracy (0.8) at $\lambda=0.8$. Regarding the $\Gamma$ parameter, Cannabis shows an overall decline in accuracy as $\Gamma$ increases. On the other hand, COBRE reaches its optimal accuracy (0.8) and AUC value (0.79) at $\Gamma=0.6$, but experiences a sharp performance drop at $\Gamma=0.8$. These findings suggest that Cannabis-related brain network structures are more effectively captured at lower $\lambda$ values, whereas COBRE-related brain functional abnormalities are more efficiently identified at higher $\lambda$ and moderate $\Gamma$ values.
Figure 8: Evaluation of brain network classification performance based on model parameters $\lambda$ and $\Gamma$.
Environment Setting
In our experiments, we follow the environment setting:
Table 4: Experiment Environment Details
Component Details
GPU NVIDIA GeForce RTX 3070
Python Version 3.7.6
Numpy Version 1.21.6
Pandas Version 1.3.5
Torch Version 1.13.1
Operating System Linux
Processor x86 64
Architecture 64-bit
Logical CPU Cores 32
Physical CPU Cores 24
Total Memory (GB) 62.44
Appendix I Biological Interpretation of k-hop Convergence
A faster decay in the spectral norm implies that higher-order information becomes less influential, indicating a more localized brain interaction pattern, while a slower decay suggests that longer-range dependencies are preserved in the FC network. As the k-value increases, more distant neural connections are considered, which may reflect the integration process of long-distance FC in the brain. Thus, the spectral norm convergence analysis provides quantitative insights into the scale and extent of connectivity alterations, complementing the interpretability of our tree-structured model.
Figure 9: The visualization of different decay meanings in spectral norm convergence. The left side of the curve (blue section) demonstrates faster decay, which leads to more localized brain interaction patterns, visualized by a compact, hierarchical network structure with blue nodes. The right side (orange section) shows slower decay, which preserves longer-range dependencies, illustrated by a more distributed network with orange nodes connected across greater distances.
Appendix J Age-Related Brain Network Changes in Addiction and Schizophrenia
In this section, we identified age-related FC subnetwork changes associated with mental disorders, as shown in Fig. 10. his study reveals distinct patterns of brain network connectivity across different age groups in individuals with addiction and schizophrenia through FC analysis. In addiction patients, age emerges as a critical variable significantly influencing brain network organization. Young addiction patients (18-25 years) demonstrate strong connectivity between the DMN, VN, and FPN, which may reflect the collective influence of these critical networks during the early stages of addiction development. As age increases, middle-aged addiction patients (25-40 years) show connectivity extending to the DAN, indicating enhanced attentional control mechanisms during persistent addiction. Notably, older addiction patients (40-55 years) predominantly exhibit Somatomotor network connectivity, possibly representing specific effects of long-term addiction on motor control systems. In contrast, schizophrenia patients display different age-related network change trajectories. Young schizophrenia patients (18-25 years) are primarily characterized by strong connections between Somatomotor and VN, which may relate to abnormal sensory-motor integration in early schizophrenia. Middle-aged schizophrenia patients (25-40 years) demonstrate stronger FPN activity, suggesting the importance of higher-order cognitive control networks in disease progression. Older schizophrenia patients (40-66 years) are mainly characterized by enhanced connectivity between DAN and VN, potentially reflecting changes in attentional processing and visual integration mechanisms in late-stage disease. From our research findings, we demonstrate significant heterogeneity across different age groups in various mental disorders, particularly in the progression patterns of age-dependent network changes.
Figure 10: Predicted FC changes in different age groups and affiliated functional subnetworks.
Appendix K Effects of Age and CMFC loss on Brain Tree Structure
According to Fig. 11, without parameter age and CMFC loss, the tree branches appear disorganized and fragmented, with paths lacking anatomical continuity and interpretability. However, with parameter that dynamically regulates FC based on individual age, it presents more coherent branches and clearer hierarchical structure.
Figure 11: Constructed brain Tree based on w/o $\theta$ and w/o CMFC loss. We present different modeling approaches for brain connectivity networks and their robust in brain tree construct. (A) The left panel illustrates brain connectivity network models without age parameters, while the right panel demonstrates models incorporating age parameters. (B) The left panel displays brain connectivity network models without CMFS loss function, whereas the right panel shows models utilizing CMFC loss function.
Appendix L Dataset
We presents the summary statistic of two public fMRI dataset below:
Table 5: Summary statistics of demographics in Cannabis and COBRE
Dataset Sample size Age (Mean ± SD)
Cannabis HC (n=128) 30.06 $\pm$ 10.47
Cannabis (n=195) 27.23 $\pm$ 7.73
COBRE HC (n=72) 38.31 $\pm$ 12.21
Schizophrenia (n=70) 36.04 $\pm$ 13.02
L.1 fMRI preprocessing in cannabis dataset
fMRI Data Processing Pipeline
(Kulkarni et al., 2023) The fMRI data underwent a comprehensive preprocessing pipeline. Raw data was acquired using 3T MRI scanners (Siemens Trio or Philips) with T2*-weighted and high-resolution T1-weighted structural images. The preprocessing, implemented through FMRIPREP, included intensity non-uniformity correction, skull stripping, and brain surface reconstruction using FreeSurfer. Spatial normalization to the ICBM 152 template was performed, followed by brain tissue segmentation into cerebrospinal fluid (CSF), white matter, and gray matter. The functional data underwent slice-timing correction, motion correction, and co-registration to corresponding T1w images using boundary-based registration. Physiological noise was addressed through CompCor regression. Additional denoising steps included removing participants with framewise displacement $>$ 3mm in over 5% of scan length and applying ArtRepair for remaining high-motion volumes. Combined task/nuisance regression was performed using motion parameters and aCompCor as covariates. The processed whole-brain time series were then parcellated using the Stanford functional ROIs atlas, resulting in 90 regions of interest. Quality control analyses confirmed that motion metrics did not significantly differ between groups (t-test p=0.86) and showed no correlation with classification accuracy (p=0.475).
L.2 Different Atlases According to Seven Affiliated Functional Subnetworks
Table 6: Comparison of Brain Networks between Stanford (90 ROIs) and Harvard-Oxford (118 ROIs) Atlases
Network Stanford (90 ROIs) Harvard-Oxford (118 ROIs)
Default Mode bilateral mPFC, L/R lateral angular gyrus, R superior frontal gyrus, bilateral posterior/middle/mid-posterior cingulate, bilateral anterior thalamus, L/R parahippocampal gyrus, L/R inferior/mid-temporal cortex, L medial angular gyrus, L crus cerebellum (lang), L inferior parietal cortex, L/R middle frontal gyrus (vent_DMN), R inferior cerebellum (vent_DMN) Left/Right Frontal Pole, Superior Frontal Gyrus, Middle Temporal Gyrus (anterior/posterior), Angular Gyrus, Precuneous Cortex
Visual L/R mid occipital cortex, L/R precuneus (post_sal), bilateral medial posterior/medial precuneus, R angular gyrus - precuneus, L/R ventral precuneus, L/R fusiform gyrus, L/R mid-occipital cortex Left/Right Middle/Inferior Temporal Gyrus (temporooccipital), Lateral Occipital Cortex (superior/inferior), Cuneal Cortex, Lingual Gyrus, Temporal Occipital Fusiform Cortex, Occipital Fusiform Gyrus, Occipital Pole
Dorsal Attention L/R superior frontal gyrus, L/R superior/inferior parietal cortex, L/R precentral/fronto-opercular region, L/R inferior temporal cortex, L/R cerebellum (visuospatial), R crus cerebellum Left/Right Middle Frontal Gyrus, Superior Parietal Lobule, Supramarginal Gyrus (anterior/posterior)
Frontoparietal L/R middle frontal/dlPFC, L inferior frontal (triangular), L/R inferior parietal/angular gyrus, L inferior temporal gyrus, R/L crus cerebellum (LECN/RECN), L posterior thalamus, R middle orbito-frontal cortex, R superior medial frontal gyrus, R caudate Left/Right Inferior Frontal Gyrus (pars triangularis/opercularis), Frontal Operculum Cortex
Somatomotor L/R superior temporal/auditory, R thalamus, L/R pre/post-central gyri, bilateral supplementary motor area, L/R ventral posterior nucleus of thalamus, cerebellar vermis Left/Right Precentral Gyrus, Superior Temporal Gyrus (anterior/posterior), Postcentral Gyrus, Juxtapositional Lobule Cortex, Central Opercular Cortex
Subcortical R striatum/thalamus (2), L/R inferior frontal gyrus, pons/dropout region Left/Right Thalamus, Caudate, Putamen, Pallidum, Brain-Stem, Hippocampus, Amygdala, Accumbens
Others L prefrontal cortex, L/R anterior insula, bilateral ACC, L/R crus cerebellum (ant_sal), L/R middle frontal gyrus, L/R supramarginal cortex, L/R middle thalamus, L/R anterior cerebellum, L/R superior temporal gyrus, bilateral calcarine cortex, L LGN Left/Right Insular Cortex, Temporal Pole, Frontal Medial/Orbital Cortex, Subcallosal Cortex, Paracingulate Gyrus, Cingulate Gyrus (anterior/posterior), Parahippocampal Gyrus, Temporal Fusiform Cortex, Parietal Operculum Cortex, Planum regions, Cerebral White Matter/Cortex, Lateral Ventricle